Grade 11General Physics

A man has to go 50 m due north, 40 m due east and 20 m due south to reach a field. (a) What distance he has to walk to reach the field? (b) What is his displacement from his house

E Sai chenchukrishnareddy
9 Years agoGrade 11

1 Answer

Gaurav Sen
9 Years ago
Total distance = 50+40+20 = 110 mVertical displacement = 30Horizontal displacement = 40Total displacement =Root of { (40)^2 + (30)^2 } = 50 m